OpenStudy (anonymous):

Find S25 for 3+7+11+15

OpenStudy (anonymous):

Let \(a_1=3\), \(d=7-3=4\) and \(n=25\). Since this is an arithmetic series, we have \[S_{25}=\frac{n}{2}[2a_1+(n-1)d]=\frac{25}{2}[2(3)+(25-1)4]=\boxed{1275}\]